# The Hitchhiker’s Guide to Human Alignment with *PO

### v1

@@ Abstract
With the growing utilization of large language models (LLMs) across domains, alignment towards human preferences has become one of the most critical aspects of training models. At the forefront of state-of-the-art human alignment methods are preference optimization methods (*PO). However, prior research has often concentrated on identifying the best-performing method, typically involving a grid search over hyperparameters, which can be impractical for general practitioners. In this paper, we aim to identify the algorithm that, while being performant, is simultaneously more robust to varying hyperparameters, thereby increasing the likelihood of achieving better results. We focus on a realistic out-of-distribution (OOD) scenario that mirrors real-world applications of human alignment, offering practical insights into the strengths and weaknesses of these methods. Furthermore, to better understand the shortcomings of generations from the different methods, we analyze the model generations through the lens of KL divergence of the SFT model and the response length statistics. Our analysis reveals that the widely adopted DPO method consistently produces lengthy responses of inferior quality that are very close to the SFT responses. Motivated by these findings, we propose an embarrassingly simple extension to the DPO algorithm, LN-DPO, resulting in more concise responses without sacrificing quality compared to the policy obtained by vanilla DPO.
The Hitchhiker’s Guide to Human Alignment with *PO
@@
Vishrav Chaudhary2                        Alon Benhaim2                        Jay Pujara1                        Xia Song2
@@ 1 Introduction
In recent years, the quality of large language models (LLMs) has been constantly increasing Chiang et al. (2024), achieving impressive results across tasks and benchmarks Abdin et al. (2024); AI@Meta (2024); Achiam et al. (2023); Team (2023); Yang et al. (2024). However, even with the most rigorous filtering heuristics, the training data Computer (2023); Penedo et al. (2024) is typically contaminated with undesirable content, that can lead to unacceptable utterances Bender et al. (2021); Gehman et al. (2020). Moreover, LLMs are pre-trained using an unsupervised next-token log-likelihood objective, which is not necessarily aligned with target behaviors.
Prior works have utilized supervised fine-tuning (SFT) as a post-training step to align LLMs with desired behaviors such as instruction-following Zhou et al. (2024); Taori et al. (2023); Chowdhery et al. (2023). To further improve the model’s alignment with human preferences, the de-facto approach has been to learn from human/AI generated preference data (e.g., a chosen and a rejected response for each prompt). In particular, off-policy preference optimization methods (*PO) have been especially popular given their good performance and ease of implementation  Rafailov et al. (2024); Hong et al. (2024); Meng et al. (2024).
@@
Mean KL Divergence 55.0 -26.0% -20.7%
Mean KL Divergence (Length-normalized) 0.5 -17.7% -3.6%
@@
One commonly occurring practice when reporting the performance of new methods is to compare their best-performing variant (after a hyperparameter grid search) to a default baseline with a fixed set of hyperparameters. However, from a practical perspective for future users, these comparisons do not provide a good answer to the problem of which method is expected to achieve higher performance, given a fixed budget for hyperparameter search, as doing broad grid searches is oftentimes computationally infeasible for many practitioners. To this end, in this work, we aim to identify the algorithm that is more robust to hyperparameter variations while still being competitive in performance. We set up our experiments in a realistic out-of-distribution (OOD) setting, focused on safety and helpfulness domains, where the train and test datasets share a common core goal, but their samples are generated from different distributions (e.g., AI and human expert). This setting resembles real-world scenarios as it simulates the release of large generative models for public use.
To better understand the behavior of vanilla DPO, we analyze the response length w.r.t. the KL and notice long responses that have low quality (i.e., very low KL). To overcome this issue, we introduce LN-DPO, which adds a simple length-regularizer (similar to SimPO) to vanilla DPO’s objective, encouraging the generation of shorter responses [^1] Concurrently, Meng et al. (2024) have added LN-DPO to their experiments (updated on July 7th, 2024). Here, we present a more thorough analysis and comparison..
In summary, our contributions are as follows:
  • We comprehensively examine DPO, LN-DPO, and SimPO across a wide range of hyperparameters in a real-world setup, providing insights into the effect of their hyperparameters on performance.
@@
  • We introduce LN-DPO, a length-normalized version of DPO that is more stable across hyperparameters, effectively reduces the average response length, and improves the performance of the trained models.
@@ 2 Related Work
Since the introduction of DPO Rafailov et al. (2024), there has been a body of works with new optimization objectives improving the performance and efficiency Azar et al. (2024); Tang et al. (2024); Hong et al. (2024); Rosset et al. (2024); Meng et al. (2024); Xu et al. (2024a); Ethayarajh et al. (2024) (see Table 2 for examples). These methods can be partitioned into two groups: reference-free Meng et al. (2024); Hong et al. (2024) and reference-dependent Rafailov et al. (2024); Park et al. (2024). Reference-free methods generally benefit from fast training runs, while reference-dependent methods have terms baked into their objective to control divergence from the reference model. In this work, we compare SimPO Meng et al. (2024), a recent state-of-the-art reference-free method, with the DPO as the reference-dependent method, providing practical insights into their performance (see Appendix A for extended related work).
@@ 3.1 Datasets
For our datasets, we follow the setup introduced by Xu et al. (2024b). Specifically, we use the double safe/unsafe filtered train subset of SafeRLHF Dai et al. (2024) for training [^2] We use the old version in our experiments: PKU-Alignment/PKU-SafeRLHF, and the test subset of HH-RLHF Ganguli et al. (2022) for evaluation. This setup closely resembles real-world scenarios where even though models are trained on various domains (e.g., safety and helpfulness in our experiments), they have to generalize to similar unseen queries while interacting with the users. For example, in our setup, both datasets focus on safety; however, SafeRLHF is human-labeled, while LLM models label HH-RLHF.
@@ 3.2 Models
For all our experiments we use the Phi-3 Medium model333microsoft/Phi-3-medium-4k-instruct Abdin et al. (2024). We chose this model due to its high performance across benchmarks and small size to ensure computational tractability. To evaluate the trained models, we use the OpenAssistant reward model [^4] OpenAssistant/oasst-rm-2.1-pythia-1.4b-epoch-2.5 Köpf et al. (2024) as an oracle to score the quality of their generated responses. We chose this model due to its small size and high performance on RewardBench Lambert et al. (2024), ensuring fast and correct evaluations for our experiments.
3.3 Metrics
@@
  • KL divergence: The summed difference of log-probabilities between the SFT and the trained models over the samples.
@@
3.4 Optimization Objectives
Considering the performances reported by Meng et al. (2024), we choose DPO as our reference-dependent method and SimPO as our reference-free method. While DPO has an implicit length normalization through the reference model, the variance of the reward (i.e., $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$) increases with response length. As such, inspired by explicit length regularization in SimPO and R-DPO Park et al. (2024), we further normalize it with the response length similar to SimPO, which we call LN-DPO. Note that this change makes LN-DPO close to an adaptive margin version of SimPO with per sample margin defined as
@@ 4.1 Hyperparameter Sensitivity
Following the common practice of comparing the best performance achieved by each method, Table 1 showcases the best result of all three models across our metrics. As evident, at their peaks, SimPO and LN-DPO beat DPO across the board. Moreover, we can observe the effectiveness of the length regularizer, drastically decreasing the mean response length (as much as 22.9%) while also having a higher performance. We also notice a significant decrease in KL divergence (as much as 26%). However, KL and length-normalized KL for SimPO decrease less than LN-DPO, showcasing a more significant divergence from SFT, which we will further analyze in the following. Overall, both SimPO and LN-DPO seem to be suitable replacements for DPO. Note that due to being reference-free, SimPO takes almost half the time to train.
@@
Given the limited resources that most users have, it is extremely difficult to run broad hyperparameter searches to find the best-performing set. As such, it becomes crucial to understand the hyperparameter sensitivity, which provides insights into the expectation of finding good hyperparameters set from a limited search. Figure 1 presents the performance distribution *PO methods following a grid search over the hyperparameters denoted in Table 2 and Appendix B. As evident, SimPO and LN-DPO effectively increase the average performance across hyperparameters, showcasing their superiority to DPO. Moreover, we observe that DPO has a relatively high rate of achieving performances worse than the SFT model, which makes the search problem more difficult. Overall, SimPO outperforms the other two models, on average and peak (i.e., the tail-end of the distribution).
Head-to-head Performance.
While comparing the pure performances achieved on the desired metrics is usually good enough to contrast different methods, there are potential cases where the averaging could be exploited (e.g., outliers with high rewards). Hence, it is essential also to do a head-to-head per sample comparison, which provides more fine-grained insights. Table 3 compares each method’s best and best 75% model (i.e., beating 75% of the other hyperparameters) based on mean score. As evident, SimPO performs best, with LN-DPO coming in second. Notably, we observe a sharp performance drop in DPO from the best to the top 25% model, in contrast to the other two. This occurrence highlights the practical flaw in only comparing the best performances.
@@ 4.2 Response Length
Since length exploitation is a critical issue  Park et al. (2024), we compare the response lengths across samples generated by the top k($k\in\{1,10,25\}$) of each method’s best-performing hyperparameters. As illustrated in Figure 2, on the best set of hyperparameters (i.e., top 1%), the non-DPO methods showcase a left shift in length distribution (compared to DPO), which is a desired effect. However, this phenomenon starts to diminish as we include worse-performing hyperparameters. For example, LN-DPO has a higher rate than DPO in the tail-end of the top 25% distribution. Overall, we observed that both length-regularized models performed superior to DPO, with SimPO producing the shortest responses across the distribution.
@@
Since reference-free methods are not regularized against a reference policy (e.g., the SFT model), there is a possibility of reward hacking occurring (i.e., lower loss with degraded performance). Therefore, we compare the KL divergence across samples generated by the top k($k\in\{1,10,25\}$) of each method’s best-performing hyperparameters. Figure 3 illustrates the KL distributions. As evident, both SimPO and LN-DPO achieve lower KLs at their peak. However, as we move toward worse-performing models, DPO achieves lower KL (at 10%). While at first, this seems like a pitfall of SimPO and LN-DPO, in reality, we believe it showcases that many DPO runs fail to learn beyond the SFT model, which is a critical shortcoming.
Since KL divergence is calculated as a summation, there is the possibility that lower values are achieved purely due to responses being shorter. As such, we also calculate a length-normalized KL to account for the effect of longer sequences, which roughly quantifies the average per token distributional difference. Notably, on this metric, SimPO always has distributionally higher values, even at the top 1% (see Figure 9, Appendix D). This observation showcases that SimPO diverges the most from the SFT on a token level, which could be explained by the lack of reference dependence.
4.4 Hyperparameter Tuning Considerations
@@
42 The Answer to the Ultimate Question
Based on our collective empirical results, we believe SimPO to be the best starting point among the three methods, mainly due to its robustness toward hyperparameter variations and effective length reduction. As for SimPO’s hyperparameters, we recommend $\beta\in\{1.0,1.5\}$ and $\gamma\approx 1.2$. Moreover, while LN-DPO is consistently second-best in most of our experiments, we discuss scenarios for choosing it over SimPO in Appendix F.
5 Conclusion
In this work, we introduce LN-DPO, a length-normalized variation of DPO that reduces the average response length while staying reference-dependent. Moreover, we present a thorough analysis of LN-DPO and two state-of-the-art reference-dependent and reference-free preference optimization methods in a simulated real-world scenario for safety and helpfulness domains. Specifically, we cover the behavior of these methods across a wide range of hyperparameters under metrics such as mean response length, KL divergence (vs. SFT), and win rate (vs. chosen and SFT). Our experiments showcase state-of-the-art methods’ strengths and weaknesses along with providing insights for other practitioners. Finally, we present our empirical insights into tuning *PO hyperparameters, potentially helping to reduce future efforts and costs.
Appendix A Extended Related Work
Online Algorithms.
Reinforcement learning from human/AI feedback (RLHF/RLAIF) is among the common approaches for aligning LLMs to human preferences Christiano et al. (2017); Bai et al. (2022a); Stiennon et al. (2020); Bai et al. (2022b), and has been used to train models such as GPT-4 Achiam et al. (2023) and Llama-3 AI@Meta (2024). In most cases, these approaches are comprised of three stages: 1) supervised fine-tuning Taori et al. (2023); Zhou et al. (2024); Xia et al. (2024), 2) reward modeling Gao et al. (2023); Chen et al. (2024); Lightman et al. (2023), and 3) policy optimization Schulman et al. (2017). The prominent method for policy optimization is Proximal Policy Optimization (PPO) which is an online on-policy approach Schulman et al. (2017). While PPO has shown promising performances Stiennon et al. (2020); Ouyang et al. (2022); Achiam et al. (2023), it suffers from problems such as having too many subtle details for reproducibility Huang et al. (2024b), 2) taking a long time for training Huang et al. (2024a), and 3) reward over-optimization Skalse et al. (2022).
Offline Algorithms.
To address the drawbacks of RLHF/RLAIF, recent works have proposed simpler and more efficient offline algorithms, particularly Direct Preference Optimization (DPO) Rafailov et al. (2024) which is based on the Bradley-Terry model Bradley and Terry (1952). These offline algorithms, directly optimize an objective on the preference data with an implicit reward model, without needing to have separate stages. Some recent works have focused on making a broad comparison between PPO and DPO. Specifically, they showcase the potential for PPO with a gold reward model ($\sim+10\%$) while underlying the similarity to DPO ($\sim+1\%$ averaged across benchmarks) when trained on the same data Ivison et al. (2024); Xu et al. (2024b).
Figure 8: Mean Length-normalized KL vs. Mean Score. Each sample presents a run with a different set of hyperparameters.
Figure 9: Length-normalized KL. The top k($k\in\{1,10,25\}$) denotes the percentage of best-performing hyperparameters taken from all the runs of each method.
Appendix B Training Regimen
Appendix C Implementation Details
Appendix D Length-Normalized KL
Figure 9 presents the length normalized KL values across samples generated by the top k($k\in\{1,10,25\}$) of each method’s best-performing hyperparameters.
Appendix E Mean Length-Normalized KL vs. Mean Score
Figure 8 presents the spread of mean length normalized KL against mean scores, across runs with different hyperparameters.
Appendix F When to use LN-DPO over SimPO?
While SimPO achieves superior performance on most metrics compared to LN-DPO, the lack of a reference policy regularization could lead to drastic divergence from the initial checkpoint, as also shown in our experiments. This issue then could cause a degradation of performance on other benchmarks, which is a critical pitfall (as also observed in Korbak et al. (2022)). As such, we believe there are various scenarios where LN-DPO should be preferred to SimPO. We leave further experiments over this direction to future works..

### v2

# A Practical Analysis of Human Alignment with *PO

@@ Abstract
At the forefront of state-of-the-art human alignment methods are preference optimization methods (*PO). Prior research has often concentrated on identifying the best-performing method, typically involving a grid search over hyperparameters, which can be impractical for general practitioners. In this paper, we examine the robustness of existing state-of-the-art methods to varying hyperparameters in a realistic out-of-distribution (OOD) scenario that mirrors real-world applications of human alignment. Our goal is to empirically find the method that increases the likelihood of achieving better results through the lens of various metrics, such as KL divergence and response length. We also introduce LN-DPO, a simple length-normalized version of DPO that is more stable across hyperparameters, effectively reduces the average response length, and improves performance. Our analysis of state-of-the-art reference-free (i.e., SimPO) and reference-dependent (i.e., DPO and LN-DPO) methods reveals that they perform similarly at their peak (i.e., best possible scenario). However, we uncover that the pattern of change in performance greatly varies as we move away from the best possible scenario.
A Practical Analysis of Human Alignment with *PO
@@
Vishrav Chaudhary2                        Alon Benhaim2                        Jay Pujara1
Xia Song2
@@ 1 Introduction
In recent years, the quality of large language models (LLMs) has been constantly increasing Chiang et al. (2024), achieving impressive results across tasks and benchmarks Abdin et al. (2024); AI@Meta (2024); Achiam et al. (2023); Team (2023); Yang et al. (2024). However, even with the most rigorous filtering heuristics, the training data Computer (2023); Penedo et al. (2024) is typically contaminated with undesirable content that can lead to unacceptable behaviors Bender et al. (2021); Gehman et al. (2020). To improve the model’s alignment with human preferences, the de-facto approach has been to learn from human/AI-generated preference data (e.g., a chosen and a rejected response for each prompt). In particular, off-policy preference optimization methods (*PO) have been prevalent given their good performance and ease of implementation  Rafailov et al. (2024); Hong et al. (2024); Meng et al. (2024).
@@
KL Divergence 55.0 -26.0% -20.7%
@@
One commonly occurring practice when reporting the performance of new methods is to compare their best-performing variant (after a hyperparameter grid search) to a default baseline with a fixed set of hyperparameters. However, from a practical perspective for future users, these comparisons do not provide a good answer to the problem of which method is expected to achieve higher performance, given a fixed budget for hyperparameter search, as doing broad grid searches is often computationally infeasible for many practitioners. To this end, in this work, we aim to empirically identify the more robust method to hyperparameter variations while still being competitive in performance.
We set up our experiments in a realistic out-of-distribution (OOD) setting, focused on safety and helpfulness domains, where the train and test datasets share a common core goal, but their samples are generated from different distributions (e.g., AI and human expert). This setting resembles real-world scenarios as it simulates the release of large generative models for public use. Moreover, to better understand the behavior of the state-of-the-art models, we take the best-performing reference-free and reference-dependent models (as reported by Meng et al. (2024)) and analyze them through the lens of standard metrics such KL divergence, response length, and win rate. We also introduce an embarrassingly simple length-normalized extension of vanilla Direct Preference Optimization (DPO) Rafailov et al. (2024), LN-DPO, that effectively mitigates the issue of lengthy generations without any apparent performance degradation [^1] Concurrently, Meng et al. (2024) have added a similar method to their experiments (updated on July 7th, 2024). Here, we present a more thorough analysis and comparison.. In summary, our contributions are as follows:
  • We examine state-of-the-art reference-free and reference-dependent preference optimization methods across a wide range of hyperparameters in a real-world setup.
@@
  • We introduce and examine LN-DPO, a simple length-normalized version of DPO that is more stable across hyperparameters, effectively reduces the average response length and improves performance.
@@ 2 Related Work
Since the introduction of DPO Rafailov et al. (2024), there has been a body of works with new optimization objectives improving the performance and efficiency Azar et al. (2024); Tang et al. (2024); Hong et al. (2024); Rosset et al. (2024); Meng et al. (2024); Xu et al. (2024a); Ethayarajh et al. (2024). These methods can be partitioned into two groups: reference-free Meng et al. (2024); Hong et al. (2024) and reference-dependent Rafailov et al. (2024); Park et al. (2024). Reference-free methods generally benefit from fast training runs, while reference-dependent methods have terms baked into their objective to control divergence from the reference model. In this work, we compare SimPO Meng et al. (2024), a recent state-of-the-art reference-free method, with DPO and LN-DPO as reference-dependent methods (see Appendix A for extended related work).
@@ 3.1 Datasets
For our datasets, we follow the setup introduced by Xu et al. (2024b). Specifically, we use the double safe/unsafe filtered train subset of SafeRLHF Dai et al. (2024) for training and the test subset of HH-RLHF Ganguli et al. (2022) for evaluation. This setup closely resembles real-world scenarios where even though models are trained on various domains (e.g., safety and helpfulness in our experiments), they have to generalize to similar unseen queries while interacting with the users.
@@ 3.2 Models
For all our experiments, we chose the Phi-3 Medium model Abdin et al. (2024) due to its high performance across benchmarks and small size, ensuring computational tractability. To evaluate the trained models, we use the OpenAssistant reward model Köpf et al. (2024) to score the quality of their generated responses. We chose this model due to its small size and use in prior works Xu et al. (2024b), ensuring fast and correct evaluations.
3.3 Optimization Objectives
Considering the performances reported by Meng et al. (2024), we choose DPO as our reference-dependent method and SimPO as our reference-free method. While DPO has an implicit length normalization through the reference model, the variance of the reward (i.e., $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$) increases with response length. As such, inspired by explicit length regularization in SimPO and R-DPO Park et al. (2024), we further normalize it with the response length similar to SimPO, which we call LN-DPO (see Section 3.4 for more details).
3.4 Connection between LN-DPO and SimPO
LN-DPO is similar to an adaptive margin version of SimPO with per sample margin defined as
4 Training Regimen
5 Metrics
@@
  • KL divergence: The summed difference of log probabilities between the SFT and the trained models over the samples.
@@
6 Implementation Details
@@ 7.1 Hyperparameter Robustness
Following the common practice, we compare the best performance achieved by each method in Table 1. As evident, at their peaks, SimPO, LN-DPO, and DPO score similarly (within a 0.05 point on average). However, SimPO and LN-DPO show an edge in terms of the rest of the metrics. Specifically, we can observe the effectiveness of the length normalization term. We also notice a significant decrease in KL divergence. However, KL for SimPO decreases less than LN-DPO, showcasing a more significant divergence from SFT. For more details on tuning these models, see Appendix B.
@@
Head-to-head Performance.
While comparing the pure performances achieved on the desired metrics is usually good enough to contrast different methods, there are potential cases where the averaging could be exploited (e.g., outliers with high rewards). Hence, it is essential also to do a head-to-head per sample comparison, which provides more fine-grained insights. Table 3 compares each method’s best and 75th percentile performance. Notably, we observe a sharp performance drop in DPO from the best to the top 25% model, in contrast to the other two. This occurrence highlights the practical flaw in only comparing the best performances.
@@
Given the limited resources that most users have, it is extremely difficult to run broad hyperparameter searches to find the best-performing combination. As such, it becomes crucial to analyze hyperparameter robustness, which provides insights into the expectation of finding good hyperparameters set from a limited search. Figure 1 presents the performance distribution *PO methods following a grid search over the hyperparameters denoted in Table 2 and Section 4. As evident, SimPO and LN-DPO effectively increase the average performance (i.e., shifting the distributions to the right) across hyperparameters, showcasing their superiority. Note that we stretched the range of hyperparameters until a plateau or an extreme variance was observed.
@@ 7.2 Response Length
Since length exploitation is a critical issue  Park et al. (2024), we compare the response lengths across samples generated by the top k($k\in\{1,10,25\}$) of each method’s best-performing hyperparameters. As illustrated in Figure 2, on the best set of hyperparameters (i.e., top 1%), the non-DPO methods showcase a left shift in length distribution (compared to DPO), which is a desired effect. However, this phenomenon starts to diminish as we include worse-performing hyperparameters. For example, LN-DPO has a higher rate than DPO in the tail-end of the top 25% distribution. Overall, we observed that both length-normalized models perform superior to DPO, with SimPO producing the shortest responses across the distribution.
@@
Since reference-free methods are not normalized against a reference policy (e.g., the SFT model), reward hacking might occur (i.e., lower loss with degraded performance). Therefore, we compare the KL divergence in Figure 3 across samples generated by the top k($k\in\{1,10,25\}$) of each method’s best-performing hyperparameters. As evident, both SimPO and LN-DPO achieve lower KLs at their peak. However, as we move toward worse-performing models, DPO achieves lower KL (at 10%). This phenomenon is due to many DPO runs failing to learn beyond the SFT model.
8 When to use LN-DPO over SimPO?
While SimPO achieves superior performance on most metrics compared to LN-DPO, the lack of a reference policy regularization could lead to drastic divergence from the initial checkpoint, as also shown in our experiments. This issue then could cause a degradation of performance on other benchmarks, which is a critical pitfall (as also observed in Korbak et al. (2022)). As such, we believe there are various scenarios where LN-DPO should be preferred to SimPO. We leave further experiments over this direction to future works.
9 Conclusion
In this work, we introduce LN-DPO, a length-normalized variation of DPO that reduces the average response length while staying reference-dependent. Moreover, we present a thorough analysis of LN-DPO and two state-of-the-art reference-dependent and reference-free preference optimization methods in a simulated real-world scenario for safety and helpfulness domains. Specifically, we cover the behavior of these methods across a wide range of hyperparameters under metrics such as mean response length, KL divergence (vs. SFT), and win rate (vs. chosen and SFT). Our experiments showcase state-of-the-art methods’ strengths and weaknesses and provide insights for other practitioners.
Limitations
Due to the extremely high costs of running such experiments (i.e., roughly 86000 GPU hours for the current experiments), in this work, we only experimented with a small set of models, methods, and datasets. While this might limit generalizability, we believe the existence of such analysis is critical to help practitioners save costs. Moreover, since the conclusion of our experiments, new reward models with higher performance have been released (e.g., ArmoRM Wang et al. (2024)); however, we still rely on older, smaller models to keep the evaluation tractable on such a high number of runs.
Acknowledgements
This work was partially funded by the Defense Advanced Research Projects Agency with the award HR00112220046.
Appendix A Extended Related Work
Online Algorithms.
Reinforcement learning from human/AI feedback (RLHF/RLAIF) is among the common approaches for aligning LLMs to human preferences Christiano et al. (2017); Bai et al. (2022a); Stiennon et al. (2020); Bai et al. (2022b), and has been used to train models such as GPT-4 Achiam et al. (2023) and Llama-3 AI@Meta (2024). In most cases, these approaches are comprised of three stages: 1) supervised fine-tuning Taori et al. (2023); Zhou et al. (2024); Xia et al. (2024), 2) reward modeling Gao et al. (2023); Chen et al. (2024); Lightman et al. (2023), and 3) policy optimization Schulman et al. (2017). The prominent method for policy optimization is Proximal Policy Optimization (PPO), an online on-policy approach Schulman et al. (2017). While PPO has shown promising performances Stiennon et al. (2020); Ouyang et al. (2022); Achiam et al. (2023), it suffers from problems such as having too many subtle details for reproducibility Huang et al. (2024b), 2) taking a long time for training Huang et al. (2024a), and 3) reward over-optimization Skalse et al. (2022).
Offline Algorithms.
To address the drawbacks of RLHF/RLAIF, recent works have proposed simpler and more efficient offline algorithms, particularly Direct Preference Optimization (DPO) Rafailov et al. (2024), which is based on the Bradley-Terry model Bradley and Terry (1952). These offline algorithms directly optimize an objective on the preference data with an implicit reward model without needing to have separate stages. Some recent works have focused on making a broad comparison between PPO and DPO. Specifically, they showcase the potential for PPO with a gold reward model ($\sim+10\%$) while underlying the similarity to DPO ($\sim+1\%$ averaged across benchmarks) when trained on the same data Ivison et al. (2024); Xu et al. (2024b).
Appendix B Hyperparameter Tuning Considerations
@@
Appendix C The Answer to the Ultimate Question
Based on our collective empirical results, we believe SimPO to be the best starting point among the three methods, mainly due to its robustness toward hyperparameter variations and effective length reduction. As for SimPO’s hyperparameters, we recommend $\beta\in\{1.0,1.5\}$ and $\gamma\approx 1.2$. Moreover, while LN-DPO is consistently second-best in most of our experiments, we discuss scenarios for choosing it over SimPO in Section 8.